# FitzHugh-Nagumo equation: from global dynamics to the slow-fast system

### v1

@@ 5. The phase portraits in the Poincaré disc
The origin is an identically zero singular point. By doing the blow-up $(u,v)\rightarrow(u_{1},w)$ where $w=\dfrac{v_{1}}{u_{1}},$ we get
$\dot{u}_{1}$ $=-\frac{a}{c}{u_{1}}^{4}{v_{1}}^{3}+\frac{1}{c}{u_{1}}^{4}{v_{1}}^{2}+c{u_{1}}^{3}{v_{1}}^{2}+\frac{b}{c}{u_{1}}^{3}{v_{1}}^{2}+c{u_{1}}^{2}{v_{1}}^{2}-\frac{c}{3}{u_{1}}^{3},$ (5.3)
$\dot{v}_{1}$ $=-c{u_{1}}^{2}{v_{1}}^{3}-cu_{1}{v_{1}}^{3}+\frac{c}{3}{u_{1}}^{2}{v_{1}}.$
By removing the common factor $u_{1},$ we have
$\dot{u}_{1}$ $=-\frac{a}{c}{u_{1}}^{3}{v_{1}}^{3}+\frac{1}{c}{u_{1}}^{3}{v_{1}}^{2}+c{u_{1}}^{2}{v_{1}}^{2}+\frac{b}{c}{u_{1}}^{2}{v_{1}}^{2}+c{u_{1}}{v_{1}}^{2}-\frac{c}{3}{u_{1}}^{2},$ (5.4)
$\dot{v}_{1}$ $=-c{u_{1}}{v_{1}}^{3}-c{v_{1}}^{3}+\frac{c}{3}{u_{1}}{v_{1}}.$
The origin is an identically zero singular point again. For studying the local phase portrait at the origin, we do another blow-up (see §2.2 of the present paper). The characteristic direction at the origin is the real factors of $\frac{2c}{3}{u_{1}}^{2}v_{1}.$ Then the vertical axis $u_{1}=0$ is a characteristic direction. So, we translate the direction $u_{1}=0$ to the direction $u_{1}=v_{1}$ doing the change of variables $(u_{1},v_{1})=(u_{2}-v_{2},v_{2}).$ Then, the new system is given by
$\dot{u_{2}}=$ $\dfrac{1}{3c}\Big{(}3c^{2}u_{2}v_{2}-c^{2}u_{2}^{2}-2c^{2}v_{2}^{2}+3c^{2}u2v_{2}^{2}+3bu_{2}^{2}v_{2}^{2}+3c^{2}u_{2}^{2}v_{2}^{2}+3u_{2}^{3}v_{2}^{2}$ (5.5)
$\qquad-6c^{2}v_{2}^{3}-6bu_{2}v_{2}^{3}-9c^{2}u_{2}v_{2}^{3}-9u_{2}^{2}v_{2}^{3}-3au_{2}^{3}v_{2}^{3}+3bv_{2}^{4}+6c^{2}v_{2}^{4}$
$\qquad+9u_{2}v_{2}^{4}+9au_{2}^{2}v_{2}^{4}-3v_{2}^{5}-9au_{2}v_{2}^{5}+3av_{2}^{6}\Big{)},$
$\dot{v_{2}}=$ $\dfrac{c}{3}v_{2}\left(u_{2}-v_{2}-3v_{2}^{2}-3u_{2}v_{2}^{2}+3v_{2}^{3}\right).$
The origin is the only singular point of system above on the line $u_{2}=0,$ and since the linear part is identically zero, we do the vertical blow-up $(u_{2},v_{2})\rightarrow(u_{3},w)$ where $w=\dfrac{v_{3}}{u_{3}},$ and we have
$\dot{u}_{3}=$ $\dfrac{1}{3c}u_{3}^{2}\Big{(}3c^{2}v_{3}-c^{2}-2c^{2}v_{3}^{2}+3c^{2}u_{3}v_{3}^{2}+3bu_{3}^{2}v_{3}^{2}+3c^{2}u_{3}^{2}v_{3}^{2}+3u_{3}^{3}v_{3}^{2}-6c^{2}u_{3}v_{3}^{3}$
$\qquad\,-6bu_{3}^{2}v_{3}^{3}-9c^{2}u_{3}^{2}v_{3}^{3}-9u_{3}^{3}v_{3}^{3}-3au_{3}^{4}v_{3}^{3}+3bu_{3}^{2}v_{3}^{4}+6c^{2}u_{3}^{2}v_{3}^{4}+9u_{3}^{3}v_{3}^{4}$
$\qquad\,+9au_{3}^{4}v_{3}^{4}-3u_{3}^{3}v_{3}^{5}-9au_{3}^{4}v_{3}^{5}+3au_{3}^{4}v_{3}^{6}\Big{)},$ (5.6)
$\dot{v}_{3}=$ $\dfrac{1}{3c}u_{3}v_{3}(1-v_{3})\Big{(}2c^{2}-2c^{2}v_{3}-6c^{2}u_{3}v_{3}^{2}-3bu_{3}^{2}v_{3}^{2}-6c^{2}u_{3}^{2}v_{3}^{2}-3u_{3}^{3}v_{3}^{2}+3bu_{3}^{2}v_{3}^{3}$
$\hskip 79.6678pt+6c^{2}u_{3}^{2}v_{3}^{3}+6u_{3}^{3}v_{3}^{3}+3au_{3}^{4}v_{3}^{3}-3u_{3}^{3}v_{3}^{4}-6au_{3}^{4}v_{3}^{4}+3au_{3}^{4}v_{3}^{5}\Big{)}.$
Eliminating the common factor $u_{3}$ leads to
$\dot{u}_{3}=$ $\dfrac{1}{3c}u_{3}\Big{(}3c^{2}v_{3}-c^{2}-2c^{2}v_{3}^{2}+3c^{2}u_{3}v_{3}^{2}+3bu_{3}^{2}v_{3}^{2}+3c^{2}u_{3}^{2}v_{3}^{2}+3u_{3}^{3}v_{3}^{2}-6c^{2}u_{3}v_{3}^{3}$
$\qquad\,-6bu_{3}^{2}v_{3}^{3}-9c^{2}u_{3}^{2}v_{3}^{3}-9u_{3}^{3}v_{3}^{3}-3au_{3}^{4}v_{3}^{3}+3bu_{3}^{2}v_{3}^{4}+6c^{2}u_{3}^{2}v_{3}^{4}+9u_{3}^{3}v_{3}^{4}$
$\qquad\,+9au_{3}^{4}v_{3}^{4}-3u_{3}^{3}v_{3}^{5}-9au_{3}^{4}v_{3}^{5}+3au_{3}^{4}v_{3}^{6}\Big{)},$ (5.7)
$\dot{v}_{3}=$ $\dfrac{1}{3c}v_{3}(1-v_{3})\Big{(}2c^{2}-2c^{2}v_{3}-6c^{2}u_{3}v_{3}^{2}-3bu_{3}^{2}v_{3}^{2}-6c^{2}u_{3}^{2}v_{3}^{2}-3u_{3}^{3}v_{3}^{2}+3bu_{3}^{2}v_{3}^{3}$
$\hskip 79.6678pt+6c^{2}u_{3}^{2}v_{3}^{3}+6u_{3}^{3}v_{3}^{3}+3au_{3}^{4}v_{3}^{3}-3u_{3}^{3}v_{3}^{4}-6au_{3}^{4}v_{3}^{4}+3au_{3}^{4}v_{3}^{5}\Big{)}.$
This system has two singular points on the line $u_{3}=0,$ namely $\tilde{E}_{1}:(0,0),$ and $\tilde{E}_{2}:(0,1).$ The eigenvalues of the Jacobian matrix associated with the vector field (5.7) at $(0,0)$ are $-\dfrac{c}{3},\dfrac{2c}{3}.$ Hence, $\tilde{E}_{1}$ is a saddle point. The linear part of the system at $\tilde{E}_{2}$ is identically zero. Therefore, in order to know the local phase portrait around this point, we must do blow-up. At first, we translate the singular point $\tilde{E}_{2}$ at the origin by doing the change of variables $(u_{3},v_{3})=(u_{4},1+v_{4}).$ System (5.7) in the new variables becomes
$\dot{u}_{4}=$ $\dfrac{1}{3c}u_{4}\Big{(}-3c^{2}u_{4}-c^{2}v_{4}-12c^{2}u_{4}v_{4}+3c^{2}u_{4}^{2}v_{4}-2c^{2}v_{4}^{2}-15c^{2}u_{4}v_{4}^{2}+3bu_{4}^{2}v_{4}^{2}$
$\qquad\quad+12c^{2}u_{4}^{2}v_{4}^{2}-6c^{2}u_{4}v_{4}^{3}+6bu_{4}^{2}v_{4}^{3}+15c^{2}u_{4}^{2}v_{4}^{3}-3u_{4}3v_{4}^{3}+3au_{4}^{4}v_{4}^{3}$
$\qquad\quad+3bu_{4}^{2}v_{4}^{4}+6c^{2}u_{4}^{2}v_{4}^{4}-6u_{4}^{3}v_{4}^{4}+9au_{4}^{4}v_{4}^{4}-3u_{4}^{3}v_{4}^{5}+9au_{4}^{4}v_{4}^{5}+3au_{4}^{4}v_{4}^{6}\Big{)},$
$\dot{u}_{4}=$ $-\dfrac{1}{3c}v_{4}(1+v_{4})\Big{(}-6c^{2}u_{4}-2c^{2}v_{4}-12c^{2}u_{4}v_{4}+3bu_{4}^{2}v_{4}+6c^{2}u_{4}^{2}v_{4}-6c^{2}u_{4}v_{4}^{2}$
$\qquad\qquad\qquad\quad+6bu_{4}^{2}v_{4}^{2}+12c^{2}u_{4}^{2}v_{4}^{2}-3u_{4}^{3}v_{4}^{2}+3au_{4}^{4}v_{4}^{2}+3bu_{4}^{2}v_{4}^{3}+6c^{2}u_{4}^{2}v_{4}^{3}$
$\hskip 85.35826pt-6u_{4}^{3}v_{4}^{3}+9au_{4}^{4}v_{4}^{3}-3u_{4}^{3}v_{4}^{4}+9au_{4}^{4}v_{4}^{4}+3au_{4}^{4}v_{4}^{5}\Big{)}.$ (5.8)
The origin is the unique singular point of system above on the line $u_{4}=0$ and furthermore, the characteristic direction at the origin of system above is $u_{4}=0.$ Hence, by using the change of variables $(u_{4},v_{4})=(u_{5}-v_{5},v_{5}),$ we translate the direction to $u_{5}=v_{5},$ and we have
$-3c\dot{u_{5}}=$ $3{u_{5}}^{4}{v_{5}}^{5}-15{u_{5}}^{3}{v_{5}}^{6}+27{u_{5}}^{2}{v_{5}}^{7}-21{u_{5}}{v_{5}}^{8}+6{v_{5}}^{9}+6{u_{5}}^{4}{v_{5}}^{4}-33{u_{5}}^{3}{v_{5}}^{5}$
$+63{u_{5}}^{2}{v_{5}}^{6}-51{u_{5}}{v_{5}}^{7}+15{v_{5}}^{8}+3{u_{5}}^{4}{v_{5}}^{3}-21{u_{5}}^{3}{v_{5}}^{4}-6c^{2}{u_{5}}^{3}{v_{5}}^{4}$
$-3b{u_{5}}^{3}{v_{5}}^{4}+12b{u_{5}}^{2}{v_{5}}^{5}+45{u_{5}}^{2}{v_{5}}^{5}+24c^{2}{u_{5}}^{2}{v_{5}}^{5}-15b{u_{5}}{v_{5}}^{6}-30c^{2}{u_{5}}{v_{5}}^{6}$
$-39{u_{5}}{v_{5}}^{6}+12c^{2}{v_{5}}^{7}+6b{v_{5}}^{7}+12{v_{5}}^{7}-3{u_{5}}^{3}{v_{5}}^{3}-15c^{2}{u_{5}}^{3}{v_{5}}^{3}-6b{u_{5}}^{3}{v_{5}}^{3}$
$+9{u_{5}}^{2}{v_{5}}^{4}+27b{u_{5}}^{2}{v_{5}}^{4}+63c^{2}{u_{5}}^{2}{v_{5}}^{4}-36b{u_{5}}{v_{5}}^{5}-9{u_{5}}{v_{5}}^{5}-81c^{2}{u_{5}}{v_{5}}^{5}$
$+33c^{2}{v_{5}}^{6}+15b{v_{5}}^{6}+3{v_{5}}^{6}-3b{u_{5}}^{3}{v_{5}}^{2}-12c^{2}{u_{5}}^{3}{v_{5}}^{2}+18b{u_{5}}^{2}{v_{5}}^{3}-27b{u_{5}}{v_{5}}^{4}$
$+60c^{2}{u_{5}}^{2}{v_{5}}^{3}-90c^{2}{u_{5}}{v_{5}}^{4}+42c^{2}{v_{5}}^{5}+12b{v_{5}}^{5}-3c^{2}{u_{5}}^{3}{v_{5}}+30c^{2}{u_{5}}^{2}{v_{5}}^{2}$
$+3b{u_{5}}^{2}{v_{5}}^{2}-69c^{2}{u_{5}}{v_{5}}^{3}-6b{u_{5}}{v_{5}}^{3}+42c^{2}{v_{5}}^{4}+3b{v_{5}}^{4}+12c^{2}{u_{5}}^{2}{v_{5}}$
$-40c^{2}{u_{5}}{v_{5}}^{2}+26c^{2}{v_{5}}^{3}+3c^{2}{u_{5}}^{2}-11c^{2}{u_{5}}{v_{5}}+6c^{2}{v_{5}}^{2},$ (5.9)
$-3c\dot{v_{5}}=$ ${v_{5}}(1+{v_{5}})(-3{u_{5}}^{3}{v_{5}}^{4}+9{u_{5}}^{2}{v_{5}}^{5}-9{u_{5}}{v_{5}}^{6}+3{v_{5}}^{7}-6{u_{5}}^{3}{v_{5}}^{3}+18{u_{5}}^{2}{v_{5}}^{4}$
$-18{u_{5}}{v_{5}}^{5}+6{v_{5}}^{6}-3{u_{5}}^{3}{v_{5}}^{2}+3b{u_{5}}^{2}{v_{5}}^{3}+6c^{2}{u_{5}}^{2}{v_{5}}^{3}+9{u_{5}}^{2}{v_{5}}^{3}-6b{u_{5}}{v_{5}}^{4}$
$-12c^{2}{u_{5}}{v_{5}}^{4}-9{u_{5}}{v_{5}}^{4}+3b{v_{5}}^{5}+6c^{2}{v_{5}}^{5}+3{v_{5}}^{5}+6b{u_{5}}^{2}{v_{5}}^{2}+12c^{2}{u_{5}}^{2}{v_{5}}^{2}$
$-24c^{2}{u_{5}}{v_{5}}^{3}-12b{u_{5}}{v_{5}}^{3}+12c^{2}{v_{5}}^{4}+6b{v_{5}}^{4}+6c^{2}{u_{5}}^{2}{v_{5}}+3b{u_{5}}^{2}{v_{5}}-18c^{2}{u_{5}}{v_{5}}^{2}$
$-6b{u_{5}}{v_{5}}^{2}+12c^{2}{v_{5}}^{3}+3b{v_{5}}^{3}-12c^{2}{u_{5}}{v_{5}}+12c^{2}{v_{5}}^{2}-6c^{2}{u_{5}}+4c^{2}{v_{5}}).$
Since the linear part is identically zero, we do the vertical blow-up $(u_{5},v_{5})=(u_{6},u_{6}v_{6}),$ for investigating the local phase portrait. The new system writes (applying the blow-up):
$-3c\dot{u}_{6}=$ ${u_{6}}^{2}(3c^{2}-11c^{2}{v_{6}}+12c^{2}{u_{6}}{v_{6}}-3c^{2}{u_{6}}^{2}{v_{6}}+6c^{2}{v_{6}}^{2}-40c^{2}{u_{6}}{v_{6}}^{2}+3b{u_{6}}^{2}{v_{6}}^{2}$
$+30c^{2}{u_{6}}^{2}{v_{6}}^{2}-3b{u_{6}}^{3}{v_{6}}^{2}-12c^{2}{u_{6}}^{3}{v_{6}}^{2}+26c^{2}{u_{6}}{v_{6}}^{3}-6b{u_{6}}^{2}{v_{6}}^{3}-69c^{2}{u_{6}}^{2}{v_{6}}^{3}$
$+18b{u_{6}}^{3}{v_{6}}^{3}+60c^{2}{u_{6}}^{3}{v_{6}}^{3}-3{u_{6}}^{4}{v_{6}}^{3}-6b{u_{6}}^{4}{v_{6}}^{3}-15c^{2}{u_{6}}^{4}{v_{6}}^{3}+3{u_{6}}^{5}{v_{6}}^{3}$
$+3b{u_{6}}^{2}{v_{6}}^{4}+42c^{2}{u_{6}}^{2}{v_{6}}^{4}-27b{u_{6}}^{3}{v_{6}}^{4}-90c^{2}{u_{6}}^{3}{v_{6}}^{4}+9{u_{6}}^{4}{v_{6}}^{4}+27b{u_{6}}^{4}{v_{6}}^{4}$
$+63c^{2}{u_{6}}^{4}{v_{6}}^{4}-21{u_{6}}^{5}{v_{6}}^{4}-3b{u_{6}}^{5}{v_{6}}^{4}-6c^{2}{u_{6}}^{5}{v_{6}}^{4}+6{u_{6}}^{6}{v_{6}}^{4}+12b{u_{6}}^{3}{v_{6}}^{5}$
$+42c^{2}{u_{6}}^{3}{v_{6}}^{5}-9{u_{6}}^{4}{v_{6}}^{5}-36b{u_{6}}^{4}{v_{6}}^{5}-81c^{2}{u_{6}}^{4}{v_{6}}^{5}+45{u_{6}}^{5}{v_{6}}^{5}+12b{u_{6}}^{5}{v_{6}}^{5}$
$+24c^{2}{u_{6}}^{5}{v_{6}}^{5}-33{u_{6}}^{6}{v_{6}}^{5}+3{u_{6}}^{7}{v_{6}}^{5}+3{u_{6}}^{4}{v_{6}}^{6}+15b{u_{6}}^{4}{v_{6}}^{6}+33c^{2}{u_{6}}^{4}{v_{6}}^{6}$
$-39{u_{6}}^{5}{v_{6}}^{6}-15b{u_{6}}^{5}{v_{6}}^{6}-30c^{2}{u_{6}}^{5}{v_{6}}^{6}+63{u_{6}}^{6}{v_{6}}^{6}-15{u_{6}}^{7}{v_{6}}^{6}+12{u_{6}}^{5}{v_{6}}^{7}$
$+6b{u_{6}}^{5}{v_{6}}^{7}+12c^{2}{u_{6}}^{5}{v_{6}}^{7}-51{u_{6}}^{6}{v_{6}}^{7}+27{u_{6}}^{7}{v_{6}}^{7}+15{u_{6}}^{6}{v_{6}}^{8}-21{u_{6}}^{7}{v_{6}}^{8}$
$+6{u_{6}}^{7}{v_{6}}^{9}),$ (5.10)
$3c\dot{v}_{6}=$ ${u_{6}}{v_{6}}({v_{6}}-1)(-9c^{2}+6c^{2}{v_{6}}-30c^{2}{u_{6}}{v_{6}}+3b{u_{6}}^{2}{v_{6}}+9c^{2}{u_{6}}^{2}{v_{6}}+26c^{2}{u_{6}}{v_{6}}^{2}$
$-6b{u_{6}}^{2}{v_{6}}^{2}-51c^{2}{u_{6}}^{2}{v_{6}}^{2}+12b{u_{6}}^{3}{v_{6}}^{2}+30c^{2}{u_{6}}^{3}{v_{6}}^{2}-3{u_{6}}^{4}{v_{6}}^{2}+3b{u_{6}}^{2}{v_{6}}^{3}$
$+42c^{2}{u_{6}}^{2}{v_{6}}^{3}-24b{u_{6}}^{3}{v_{6}}^{3}-72c^{2}{u_{6}}^{3}{v_{6}}^{3}+9{u_{6}}^{4}{v_{6}}^{3}+15b{u_{6}}^{4}{v_{6}}^{3}+33c^{2}{u_{6}}^{4}{v_{6}}^{3}$
$-12{u_{6}}^{5}{v_{6}}^{3}+12b{u_{6}}^{3}{v_{6}}^{4}+42c^{2}{u_{6}}^{3}{v_{6}}^{4}-9{u_{6}}^{4}{v_{6}}^{4}-30b{u_{6}}^{4}{v_{6}}^{4}-66c^{2}{u_{6}}^{4}{v_{6}}^{4}$
$+36{u_{6}}^{5}{v_{6}}^{4}+6b{u_{6}}^{5}{v_{6}}^{4}+12c^{2}{u_{6}}^{5}{v_{6}}^{4}-15{u_{6}}^{6}{v_{6}}^{4}+3{u_{6}}^{4}{v_{6}}^{5}+15b{u_{6}}^{4}{v_{6}}^{5}$
$+33c^{2}{u_{6}}^{4}{v_{6}}^{5}-36{u_{6}}^{5}{v_{6}}^{5}-12b{u_{6}}^{5}{v_{6}}^{5}-24c^{2}{u_{6}}^{5}{v_{6}}^{5}+45{u_{6}}^{6}{v_{6}}^{5}-6{u_{6}}^{7}{v_{6}}^{5}$
$+12{u_{6}}^{5}{v_{6}}^{6}+6b{u_{6}}^{5}{v_{6}}^{6}+12c^{2}{u_{6}}^{5}{v_{6}}^{6}-45{u_{6}}^{6}{v_{6}}^{6}+18{u_{6}}^{7}{v_{6}}^{6}+15{u_{6}}^{6}{v_{6}}^{7}$
$-18{u_{6}}^{7}{v_{6}}^{7}+6{u_{6}}^{7}{v_{6}}^{8}).$
Now, by elimination of the common factor $u_{6},$ we obtain the truncated system
$\dot{u_{6}}=$ $-\dfrac{1}{3c}u_{6}\left(3c^{2}-11c^{2}v_{6}+\mathcal{O}(|u_{6},v_{6}|^{2})\right),$ (5.11)
$\dot{v_{6}}=$ $\dfrac{1}{3c}v_{6}(v_{6}-1)\left(-9c^{2}+6c^{2}v_{6}+\mathcal{O}(|u_{6},v_{6}|^{2})\right).$
The equilibrium points for system (5.11) on the line $u_{6}=0$ are given by
$\tilde{E}_{4}=(0,0),\qquad\tilde{E}_{5}=(0,1),\qquad\tilde{E}_{6}=\left(0,\frac{3}{2}\right).$ (5.12)
The eigenvalues of Jacobian matrix associated with vector field (5.11) at $\tilde{E}_{4}$ and $\tilde{E}_{5}$ are $-c$, $3c$ and $\frac{2c}{3},-c$. So, they are saddle points. On the other hand, the eigenvalues associated with Jacobian matrix at $\tilde{E}_{6}$ are 00 and $\frac{3c}{2}$, and it is a semi-hyperbolic equilibrium. Given the center manifold $\dot{u_{6}}=-\frac{b}{8c}{u_{6}}^{3}$, we conclude that the equilibrium point $\tilde{E}_{6}$ is:

### v3

@@ 5. The phase portraits “near infinity”
The origin is an identically zero singular point for system (5.3). In order to desingularize this equilibrium, we are going to employ the quasi-homogeneous blow-up
$(u,v)=(\bar{u}\rho^{2},\bar{v}\rho^{3}),$ (5.4)
where $(\alpha,\beta)=(2,3)$ has been obtain by using the Newton diagram discussed in Subsection 2.2. The Newton polygon associated to (5.3) is illustrated in Figure 9.
Figure 9. The points/segments associated to (5.3) are $(2,2)$, $(1,3)$, $(3,0)$, $(1,2)$, $(0,2)$, $(0,4)$ and $(0,3)$. The Newton polygon is defined by the line $3y+2x\geq 6$. Therefore $(\alpha,\beta)=(2,3)$.
Applying the transformation (5.4) where $\bar{u}=1$ and $\bar{u}=-1,$ the $x$ directional blow-up, converts the system (5.3) into the vector fields defined by:
$X_{1}^{+}:\begin{cases}\frac{d{\rho}}{d\eta}=\frac{\rho\left(-3a\rho^{5}{\bar{v}}^{3}+3{\rho}^{4}{\bar{v}}^{2}+3{c}^{2}{\rho}^{2}{\bar{v}}^{2}+3b{\rho}^{2}{\bar{v}}^{2}+3{c}^{2}{\bar{v}}^{2}-{c}^{2}\right)}{6c},\\ \\ \frac{d\bar{v}}{d\eta}=-\frac{\bar{v}\left(-a\rho^{5}{\bar{v}}^{3}+{\rho}^{4}{\bar{v}}^{2}+3{c}^{2}{\rho}^{2}{\bar{v}}^{2}+b{\rho}^{2}{\bar{v}}^{2}+3{c}^{2}{\bar{v}}^{2}-{c}^{2}\right)}{2c},\end{cases}$
and
$X_{1}^{-}:\begin{cases}\frac{d{\rho}}{d\eta}=\frac{\rho\left(-3a\rho^{5}{\bar{v}}^{3}-3{\rho}^{4}{\bar{v}}^{2}+3{c}^{2}{\rho}^{2}{\bar{v}}^{2}+3b{\rho}^{2}{\bar{v}}^{2}-3{c}^{2}{\bar{v}}^{2}-{c}^{2}\right)}{6c},\\ \\ \frac{d\bar{v}}{d\eta}=-\frac{\bar{v}\left(-a\rho^{5}{\bar{v}}^{3}+5{\rho}^{4}{\bar{v}}^{2}+3{c}^{2}{\rho}^{2}{\bar{v}}^{2}+b{\rho}^{2}{\bar{v}}^{2}+3{c}^{2}{\bar{v}}^{2}-{c}^{2}\right)}{2c},\end{cases}$
respectively, where $d\eta=\rho^{4}d\tau$ (the old time variable is $\tau$ and the new one is $\eta$; when we calculate $X^{1}$, we can cancel $\rho^{4}$ by using this time rescaling).
We now need to analyze the equilibria at $\rho=0$ of $X_{1}^{\pm}.$ Regarding the vector field $X_{1}^{+},$ and $X_{1}^{-},$ one concludes that:
  1. Claim 1:
    The equilibria of system $X_{1}^{+}$ on $\rho=0$ are directly determined by the roots of
    $\bar{v}\left(3\bar{v}^{2}-1\right)=0.$
    There are three equilibria at $\rho=0$ denoted by
    $D_{1}:(0,0)\qquad\text{and}\qquad f^{\pm}:\left(0,\pm\dfrac{\sqrt{3}}{3}\right).$
    The following statements are valid: